OpenStudy (anonymous):

3x+y= -7 x-y=-5 Explain whether to use elimination or substitution

OpenStudy (dwelsher):

for me it is easier to use substitution. On the bottom equation y=x+5 and therefore substitute 3x+(x+5)=-5 and then solve for x.( the only reason I am using this is because I forgot the elimination method!!

OpenStudy (gudden):

@you can use any method for solving the equations... here you go with substitution method... ( it is a method in which we substitute the value of a variable from one equation to that in the other) from first equation the value of y comes to be 3x+y= -7 y = -7 - 3x now substituting it in second equation you can get the values of x and then on keeping the value of x in any of the equation u could get the value of y and elimination method means to eliminate either either of the x or y and here as you can see if you subtract second equation from the first one you could eliminate the "y" variable... and then keeping the value of it in any of the equations you could get the other variable too..!!

OpenStudy (gudden):

so... using substitution or elimination depends on u... if u r more comfortable in solving using Elimination then use it or if u r more comfortable in using Substitution then use it...:D

OpenStudy (anonymous):

@gudden thank you so much for the great explanation! ;D